OpenStudy (anonymous):

What are the polar coordinates for the point (1,-2) i got r = 5 but theta = arctan -2?

OpenStudy (anonymous):

\[r=\sqrt{5}\]

OpenStudy (anonymous):

yeah i got that far

OpenStudy (anonymous):

and yes, in this case \[\theta =\tan^{-1}(-2)\] because you are in quadrant 4

OpenStudy (anonymous):

okay so arctan(-2) is correct?

OpenStudy (anonymous):

yes. it is always the case that \[\tan(\theta)=\frac{b}{a}\] but it is not always the case that \[\theta=\tan^{-1}(\frac{b}{a})\]

OpenStudy (anonymous):

because arctan only returns answers in \[[-\frac{\pi}{2},\frac{\pi}{2}]\]so you need to be careful about what quadrant you are in